# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

### v1

@@ C.2 Adaptivity of the Robust Confidence Set
Assume that there exist $A_{0},B_{0},C_{0},E_{0},F_{0}\in\mathbb{R}$ with $B_{0}\neq 0$ such that $A_{N}=A_{0}+o_{P}(1)$, $B_{N}=B_{0}+o_{P}(1)$, $C_{N}=C_{0}+o_{P}(1)$, $E_{N}=E_{0}+o_{P}(1)$ and $F_{N}=F_{0}+o_{P}(1)$. Then, the confidence set $\mathcal{C}_{N}^{\alpha}(a)$ is asymptotically equivalent to the standard confidence interval, i.e. for fixed $\alpha\in(0,1)$,
$C_{N}^{\alpha}(a)=\left[\hat{\beta}(a)-\frac{z_{1-\alpha/2}\hat{\sigma}(a)}{\sqrt{Nh}}+o_{P}\left(\frac{1}{\sqrt{Nh}}\right),\,\hat{\beta}(a)+\frac{z_{1-\alpha/2}\hat{\sigma}(a)}{\sqrt{Nh}}+o_{P}\left(\frac{1}{\sqrt{Nh}}\right)\right]$
@@
$\mathcal{C}_{N}^{\alpha}(a)=\left[\frac{A_{N}B_{N}\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{B_{N}^{2}C_{N}+A_{N}^{2}F_{N}-A_{N}B_{N}E_{N}+O_{P}\left(\frac{1}{Nh}\right)}+O_{P}\left(\frac{1}{Nh}\right)}{B_{N}^{2}+O_{P}\left(\frac{1}{Nh}\right)}\right].$
@@
$\textstyle\mathcal{C}_{N}^{\alpha}(a)$ $=\textstyle\left[\frac{A_{N}}{B_{N}}\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{\frac{C_{N}}{B_{N}^{2}}+\frac{A_{N}^{2}F_{N}}{B_{N}^{4}}-\frac{A_{N}E_{N}}{B_{N}^{3}}+O_{P}\left(\frac{1}{Nh}\right)}+O_{P}\left(\frac{1}{Nh}\right)\right]$
$=\textstyle\left[\hat{\beta}(a)\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{\frac{C_{N}}{B_{N}^{2}}+\hat{\beta}(a)^{2}\frac{F_{N}}{B_{N}^{2}}-\hat{\beta}(a)\frac{E_{N}}{B_{N}^{2}}}+O_{P}(1/(Nh))\right]$
where by (23), $\hat{\beta}(a)=A_{N}/B_{N}$. Note that
$\frac{C_{N}}{B_{N}^{2}}+\hat{\beta}(a)^{2}\frac{F_{N}}{B_{N}^{2}}-\hat{\beta}(a)\frac{E_{N}}{B_{N}^{2}}=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(R_{Y,i}^{k}-\hat{\beta}(a)R_{D,i}^{k})^{2}(R_{f,i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}}{\left(\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right)^{2}},$
which is equal to $\hat{\sigma}^{2}(a)$ defined in (24). Hence, the robust and the non-robust confidence intervals are asymptotically equivalent. ∎
@@ D.1 Proof of Theorem 12
The strategy is to replace the estimated nuisance functions $\hat{f}^{k},\hat{\phi}^{k},\hat{l}^{k}$ by the true functions $f,\phi,l$ in $\hat{Q}(\beta_{0}(a),a)$. In fact, the theorem follows from the following two lemmas.
@@
$\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(a)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(a)(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{A_{i}-a}{h}\right)$
$\to\mathcal{N}(0,1),\text{ in distribution.}$.
@@
$\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(a)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(a)(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{A_{i}-a}{h}\right)$
$\textstyle-\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(a),a)}{\sigma_{Q}(a)}=o_{P}(1).$
@@ D.1.1 Proof of Lemma 18
Under $H_{0}$, we have $\beta(a)=\beta_{0}(a)$ and hence $Y_{i}-l(X_{i})-\beta_{0}(a)(D_{i}-\phi(X_{i}))=\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))$. Hence, we need to consider the i.i.d. random variables
$W_{i}=\frac{\left(\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{A_{i}-a}{h}\right)}{\sqrt{h}\sigma_{Q}(a)}.$
By Assumption 2, $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$. We can use the tower property of conditional expectations together with the fact that $A_{i}$ is measurable with respect to $X_{i}$, $\mathbb{E}[D_{i}|Z_{i},X_{i}]=f(Z_{i},X_{i})$ and Lemma 24 from Section D.5 below to arrive at
$\mathbb{E}[W_{i}]$ $=\frac{1}{\sqrt{h}\sigma_{Q}(a)}\mathbb{E}\left[(\beta(A_{i})-\beta(a))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}]K\left(\frac{A_{i}-a}{h}\right)\right]$
$=h^{2.5}\frac{\rho^{\prime\prime}(a)}{2\sigma_{Q}(a)}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x+\frac{R(a,h)}{\sqrt{h}\sigma_{Q}(a)}.$
with $\rho(x)=(\beta(x)-\beta(a))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ and $R(a,h)$ defined accordingly in Lemma 24. By assertion 4 of Assumption 6, there exists a constant $C\in\mathbb{R}$ independent of $N$ such that
@@
$\textstyle\mathbb{E}[W_{i}^{2}]=\textstyle\frac{1}{h\sigma_{Q}^{2}(a)}\mathbb{E}\left[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}\right]$
$\quad\textstyle+\frac{2}{h\sigma_{Q}^{2}(a)}\mathbb{E}\left[\epsilon_{i}(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}\right]$
$\quad\textstyle+\frac{1}{h\sigma_{Q}^{2}(a)}\mathbb{E}\left[(\beta(A_{i})-\beta(a))^{2}(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}\right]$
$\textstyle=\frac{1}{\sigma_{Q}^{2}(a)}\mathbb{E}\left[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a\right]p_{A}(a)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+O(h^{2})$
@@
by the definition of $\sigma_{Q}^{2}(a)$. The quantity from Lemma 18 is $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{i}$. We apply the Lindeberg-Feller Theorem (see for example [15], Thm. 3.4.10.) to the random variables $(W_{i}-\mathbb{E}[W_{i}])/\sqrt{N}$. For this, note that
@@
$\textstyle=2^{3+\eta}\frac{\mathbb{E}\left[\left|\left(\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{A_{i}-a}{h}\right)\right|^{2+\eta}\right]}{{\delta^{\eta}N^{\eta/2}h^{\eta/2}\sigma_{Q}(a)^{2+\eta}}\cdot h}.$
This converges to 00 for $N\to\infty$, because $Nh\to\infty$ and because of assertion 12 of Assumption 6. From the Lindeberg-Feller Theorem, we can conclude that
@@ D.1.2 Proof of Lemma 19
$\Delta l_{i}^{k}=l(X_{i})-\hat{l}^{k}(X_{i}),\quad\Delta f_{i}^{k}=f(Z_{i},X_{i})-\hat{f}^{k}(Z_{i},X_{i}),\quad\Delta\phi_{i}^{k}=\phi(X_{i})-\hat{\phi}^{k}(X_{i}).$
Under $H_{0}:\beta(a)=\beta_{0}(a)$ we can write (using that $Y_{i}-l(X_{i})-\beta(a)(D_{i}-\phi(X_{i}))=\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))$ in the last step)
$\textstyle\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(a),a)}{\sigma_{Q}(a)}$
$=\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(a)(D_{i}-\hat{\phi}^{k}(X_{i})))\cdot$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))K\left(\frac{A_{i}-a}{h}\right)$
$=\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-l(X_{i})-\beta(a)(D_{i}-\phi(X_{i}))+\Delta l_{i}^{k}-\beta(a)\Delta\phi_{i}^{k}\right)\cdot$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad\cdot\left(f(Z_{i},X_{i})-\phi(X_{i})-\Delta f_{i}^{k}+\Delta\phi_{i}^{k}\right)K\left(\frac{A_{i}-a}{h}\right)$
$\textstyle=\frac{1}{\sqrt{Nh}\sigma_{Q}(a)}\sum_{i=1}^{N}\left(Y_{i}-l(X_{i})-\beta(a)(D_{i}-\phi(X_{i}))\right)\cdot$
$\textstyle\quad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))K\left(\frac{A_{i}-a}{h}\right)$
$\textstyle+\frac{1}{\sqrt{Nh}\sigma_{Q}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\Bigl{[}\left(\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\right)\left(\Delta\phi_{i}^{k}-\Delta f_{i}^{k}\right)$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(a)\Delta\phi_{i}^{k}\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(a)\Delta\phi_{i}^{k}\right)\left(\Delta\phi_{i}^{k}-\Delta f_{i}^{k}\right)\Bigr{]}K\left(\frac{A_{i}-a}{h}\right).$ (30)
@@
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (31)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\epsilon_{i}\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (32)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (33)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (34)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (35)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (36)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (37)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (38)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\left(\Delta\phi_{i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1)$ (39)
$\frac{1}{\sqrt{nh}\sigma_{Q}(a)}\sum_{i\in I_{k}}\Delta\phi_{i}^{k}\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1).$ (40)

### v2

@@ C.2 Adaptivity of the Robust Confidence Set
Assume that there exist $A_{0},B_{0},C_{0},E_{0},F_{0}\in\mathbb{R}$ with $B_{0}\neq 0$ such that $A_{N}=A_{0}+o_{P}(1)$, $B_{N}=B_{0}+o_{P}(1)$, $C_{N}=C_{0}+o_{P}(1)$, $E_{N}=E_{0}+o_{P}(1)$ and $F_{N}=F_{0}+o_{P}(1)$. Then, the confidence set $\mathcal{C}_{N}^{\alpha}(v)$ is asymptotically equivalent to the standard confidence interval, i.e. for fixed $\alpha\in(0,1)$,
$C_{N}^{\alpha}(v)=\left[\hat{\beta}(v)-\frac{z_{1-\alpha/2}\hat{\sigma}(v)}{\sqrt{Nh}}+o_{P}\left(\frac{1}{\sqrt{Nh}}\right),\,\hat{\beta}(v)+\frac{z_{1-\alpha/2}\hat{\sigma}(v)}{\sqrt{Nh}}+o_{P}\left(\frac{1}{\sqrt{Nh}}\right)\right]$
@@
$\mathcal{C}_{N}^{\alpha}(v)=\left[\frac{A_{N}B_{N}\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{B_{N}^{2}C_{N}+A_{N}^{2}F_{N}-A_{N}B_{N}E_{N}+O_{P}\left(\frac{1}{Nh}\right)}+O_{P}\left(\frac{1}{Nh}\right)}{B_{N}^{2}+O_{P}\left(\frac{1}{Nh}\right)}\right].$
@@
$\textstyle\mathcal{C}_{N}^{\alpha}(v)$ $=\textstyle\left[\frac{A_{N}}{B_{N}}\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{\frac{C_{N}}{B_{N}^{2}}+\frac{A_{N}^{2}F_{N}}{B_{N}^{4}}-\frac{A_{N}E_{N}}{B_{N}^{3}}+O_{P}\left(\frac{1}{Nh}\right)}+O_{P}\left(\frac{1}{Nh}\right)\right]$
$=\textstyle\left[\hat{\beta}(v)\pm\frac{z_{1-\alpha/2}}{\sqrt{Nh}}\sqrt{\frac{C_{N}}{B_{N}^{2}}+\hat{\beta}(v)^{2}\frac{F_{N}}{B_{N}^{2}}-\hat{\beta}(v)\frac{E_{N}}{B_{N}^{2}}}+O_{P}(1/(Nh))\right]$
where by (23), $\hat{\beta}(v)=A_{N}/B_{N}$. Note that
$\frac{C_{N}}{B_{N}^{2}}+\hat{\beta}(v)^{2}\frac{F_{N}}{B_{N}^{2}}-\hat{\beta}(v)\frac{E_{N}}{B_{N}^{2}}=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(R_{Y,i}^{k}-\hat{\beta}(v)R_{D,i}^{k})^{2}(R_{f,i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}}{\left(\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}},$
which is equal to $\hat{\sigma}^{2}(v)$ defined in (24). Hence, the robust and the non-robust confidence intervals are asymptotically equivalent. ∎
@@ D.1 Proof of Theorem 11
The strategy is to replace the estimated nuisance functions $\hat{f}^{k},\hat{\phi}_{1}^{k},\hat{\phi}_{2}^{k},\hat{l}^{k}$ by the true functions $f,\phi,l$ in $\hat{Q}(\beta_{0}(v),v)$. In fact, the theorem follows from the following two lemmas.
@@
$\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$\to\mathcal{N}(0,1),\text{ in distribution.}$
@@
$\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{i=1}^{N}(Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i})))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$\textstyle-\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(v),v)}{\sigma_{Q}(v)}=o_{P}(1).$
@@ D.1.1 Proof of Lemma 17
Under $H_{0}$, we have $\beta(v)=\beta_{0}(v)$ and hence $Y_{i}-l(X_{i})-\beta_{0}(v)(D_{i}-\phi(X_{i}))=\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))$. Hence, we need to consider the i.i.d. random variables
$W_{i}=\frac{\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right)}{\sqrt{h}\sigma_{Q}(v)}.$
By Assumption 2, $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$. We can use the tower property of conditional expectations together with the fact that $V_{i}$ is measurable with respect to $X_{i}$, $\mathbb{E}[D_{i}|Z_{i},X_{i}]=f(Z_{i},X_{i})$ and Lemma 23 from Section D.5 below to arrive at
$\mathbb{E}[W_{i}]$ $=\frac{1}{\sqrt{h}\sigma_{Q}(v)}\mathbb{E}\left[(\beta(V_{i})-\beta(v))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}]K\left(\frac{V_{i}-v}{h}\right)\right]$
$=h^{2.5}\frac{\rho^{\prime\prime}(v)}{2\sigma_{Q}(v)}\int_{-\infty}^{\infty}K(x)x^{2}\mathrm{d}x+\frac{R(v,h)}{\sqrt{h}\sigma_{Q}(v)}.$
with $\rho(x)=(\beta(x)-\beta(v))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ and $R(v,h)$ defined accordingly in Lemma 23. By assertion 4 of Assumption 6, there exists a constant $C\in\mathbb{R}$ independent of $N$ such that
@@
$\textstyle\mathbb{E}[W_{i}^{2}]=\textstyle\frac{1}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]$
$\quad\textstyle+\frac{2}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[\epsilon_{i}(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]$
$\quad\textstyle+\frac{1}{h\sigma_{Q}^{2}(v)}\mathbb{E}\left[(\beta(V_{i})-\beta(v))^{2}(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]$
$\textstyle=\frac{1}{\sigma_{Q}^{2}(v)}\mathbb{E}\left[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v\right]p_{V}(v)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+O(h^{2})$
@@
by the definition of $\sigma_{Q}^{2}(v)$. The quantity from Lemma 17 is $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{i}$. We apply the Lindeberg-Feller Theorem (see for example [20], Thm. 3.4.10.) to the random variables $(W_{i}-\mathbb{E}[W_{i}])/\sqrt{N}$. For this, note that
@@
$\textstyle=2^{3+\eta}\frac{\mathbb{E}\left[\left|\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right)\right|^{2+\eta}\right]}{{\delta^{\eta}N^{\eta/2}h^{\eta/2}\sigma_{Q}(v)^{2+\eta}}\cdot h}.$
This converges to 0 for $N\to\infty$, because $Nh\to\infty$ and because of assertion 12 of Assumption 6. From the Lindeberg-Feller Theorem, we can conclude that
@@ D.1.2 Proof of Lemma 18
$\Delta l_{i}^{k}$ $=l(X_{i})-\hat{l}^{k}(X_{i}),$ $\quad\Delta f_{i}^{k}$ $=f(Z_{i},X_{i})-\hat{f}^{k}(Z_{i},X_{i}),$
$\Delta\phi_{1,i}^{k}$ $=\phi(X_{i})-\hat{\phi}_{1}^{k}(X_{i}),$ $\quad\Delta\phi_{2,i}^{k}$ $=\phi(X_{i})-\hat{\phi}_{2}^{k}(X_{i})$
Under $H_{0}:\beta(v)=\beta_{0}(v)$ we can write (using that $Y_{i}-l(X_{i})-\beta(v)(D_{i}-\phi(X_{i}))=\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))$ in the last step)
$\textstyle\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(v),v)}{\sigma_{Q}(v)}$
$=\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i})))\cdot$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad\cdot(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$=\textstyle\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-l(X_{i})-\beta(v)(D_{i}-\phi(X_{i}))+\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\cdot$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad\cdot\left(f(Z_{i},X_{i})-\phi(X_{i})-\Delta f_{i}^{k}+\Delta\phi_{2,i}^{k}\right)K\left(\frac{V_{i}-v}{h}\right)$
$\textstyle=\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{i=1}^{N}\left(Y_{i}-l(X_{i})-\beta(v)(D_{i}-\phi(X_{i}))\right)\cdot$
$\textstyle\quad\qquad\qquad\qquad\quad\cdot(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$\textstyle+\frac{1}{\sqrt{Nh}\sigma_{Q}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\Bigl[\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k}\right)$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)$
$\textstyle\qquad\qquad\qquad\qquad\qquad\quad+\left(\Delta l_{i}^{k}-\beta(v)\Delta\phi_{1,i}^{k}\right)\left(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k}\right)\Bigr]K\left(\frac{V_{i}-v}{h}\right).$ (30)
@@
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\epsilon_{i}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (31)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\epsilon_{i}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (32)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (33)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\Delta\phi_{i,2}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (34)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta l_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (35)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{i,1}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (36)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (37)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta l_{i}^{k}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (38)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta\phi_{1,i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1)$ (39)
$\frac{1}{\sqrt{nh}\sigma_{Q}(v)}\sum_{i\in I_{k}}\Delta\phi_{1,i}^{k}\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1).$ (40)